OpenStudy (anonymous):

What is the solution to the rational equation 3/2 - 1/x-4 = 1/ 2x-8 ?

OpenStudy (pfenn1):

\[\left( 3 \over 2 \right)-\left(\begin{matrix}1 \over x-4 \\\end{matrix}\right)=\left( 1 \over 2x-8 \right)\] Multiply both sides by 2(x-4) and simplify

OpenStudy (anonymous):

3,2,5, or 4?

OpenStudy (pfenn1):

You mean what is the answer? 3,2,5,or 4? Well, what answer did you get?

OpenStudy (anonymous):

Yeah, which one is the answer? I got 2

OpenStudy (pfenn1):

Recheck your math because 2 is not the answer.

OpenStudy (anonymous):

Can you just tell me the answer? Last question on my homework. Please

OpenStudy (pfenn1):

I multiplied each side by 2 (x-4) and got\[3(x-4)-2=1\]\[3x-14=1\]\[3x=15\]\[x=5\]

OpenStudy (anonymous):

Your a sweetheart

OpenStudy (pfenn1):

You are welcome